# Orbifold modifications of complex analytic varieties

### v4

@@ 2.C. Notion of complex analytic orbispaces
    Let $U$ be a connected Hausdorff space. An orbifold chart of $U$ is a triple $(\widetilde{U},G,\pi)$ such that $\widetilde{U}$ is a reduced complex analytic space of dimension $n$, that $G$ is a finite group acting holomorphically on $\widetilde{U}$, and that $\pi\colon\widetilde{U}\to U$ is a continuous map inducing a homeomorphism from $\widetilde{U}/G$ to $U$. We denote by $\mathrm{ker}\,G$ the maximal subgroup of $G$ which acts trivially on $\widetilde{U}$.
@@
    Let $\iota\colon U^{\prime}\hookrightarrow U$ be an open embedding, and let $(\widetilde{U}^{\prime},G^{\prime},\pi^{\prime})$ be an orbifold chart of $U^{\prime}$. An injection from $(\widetilde{U}^{\prime},G^{\prime},\pi^{\prime})$ to $(\widetilde{U},G,\pi)$ consists of an open embedding $\varphi\colon\widetilde{U}^{\prime}\to\widetilde{U}$ and a group monomorphism $\lambda\colon G^{\prime}\to G$ such that $\varphi$ is equivariant with respect to $\lambda$, that $\lambda$ induces an isomorphism from $\mathrm{ker}\,G^{\prime}$ to $\mathrm{ker}\,G$, and that $\iota\circ\pi^{\prime}=\pi\circ\varphi$.
@@
Assume that $X$ is a complex analytic variety with klt singularities. Then for every point $x\in X$, there is an open neighborhood $U(x)$, and a smallest positive integer $k(x)$, such that $(\omega_{X}^{\otimes k(x)})^{**}|_{U(x)}\cong\mathcal{O}_{U(x)}$, where $\omega_{X}$ is the canonical sheaf of $X$. Let $V(x)\to U(x)$ be the cyclic cover induced by this isomorphism (see [KM98, Definition 2.52]), and let $G(x)$ be the Galois group. Then $V(x)$ has canonical singularities. Furthermore, the collection of the $(V(x),G(x))$’s induces a complex analytic orbispace with quotient space $X$, see [DO23, Lemma 2.3].
@@
Let $\mathfrak{X}=(X_{i},G_{i})_{i\in I}$ be a complex analytic orbispace with quotient space $X$. Let $\delta_{i}$ be a holomorphic function on $X_{i}$ such that the ideal sheaf $\mathcal{J}_{i}$ generated by $\delta_{i}$ is compatible along the overlaps. In other words, if $p_{i}\colon V\to X_{i}$ and $p_{j}\colon V\to X_{j}$ are inclusions of orbifold charts, then $p_{i}^{*}\mathcal{J}_{i}=p_{j}^{*}\mathcal{J}_{j}$ are the same ideal sheaf on $V$. Let $S\subseteq X$ be a compact subset and let $S_{i}=\pi_{i}^{-1}(S)$.
Then there is a finite family $(U_{k},G^{\prime}_{k})_{k\in K}$ of orbifold charts satisfying the following properties. There is an application $\sigma\colon K\to I$ and there are inclusions $\iota_{k}\colon U_{k}\to X_{\sigma(k)}$ of orbifold charts. Let $n>0$ be an integer and let $W_{k}=U_{k}[\sqrt[n]{\delta_{\sigma(k)}}]$. Then the natural morphism $W_{k}\to X$ is Galois over its image with Galois group $H_{k}$, and the family $(W_{k},H_{k})_{k\in K}$ induces a complex analytic orbispace, whose quotient space is an open neighborhood of $S$.
@@
For any $i\in I$, the assumption implies that there is a unit function $u_{g}$ such that $g^{*}\delta_{i}=u_{g}\cdot\delta_{i}$ for any $g\in G_{i}$. For every point $x\in S_{i}$, there is an open neighborhood $U_{x}$ of $x$ in $X_{i}$, such that $u_{g}|_{U_{x}}$ admits $n$-th roots. Let $G_{x}\subseteq G_{i}$ be the stabilizer of $x$. By shrinking $U_{x}$, we may assume that $(U_{x},G_{x})$ is an orbifold chart compatible with $\mathfrak{X}$. By the same argument of Lemma 2.5, the automorphisms of $U_{x}$ over $U_{x}/G_{x}$ can be lifted to automorphisms of $U_{x}[\sqrt[n]{\delta_{i}}]$. This implies that the natural morphism $U_{x}[\sqrt[n]{\delta_{i}}]\to X$ is Galois over its image.
By considering all points $x\in S_{i}$ and all $i\in I$, we obtain a family of orbifold charts $(U_{x},G_{x})$. Since $S$ is compact, we can extract a finite family of points $(x_{k})_{k\in K}$, with an application $\sigma\colon K\to I$, such that $x_{k}\in X_{\sigma(k)}$, and that the images of $U_{x_{k}}$ in $X$ covers $S$. We let $U_{k}=U_{x_{k}}$ and $G^{\prime}_{k}=G_{x_{k}}$. Then the previous paragraph implies that $W_{k}\to X$ is Galois over its image, with Galois group $H_{k}$. By Lemma 2.5 and Lemma 2.11, we deduce that the family $(W_{k},H_{k})_{k\in K}$ induces a complex analytic orbispace. ∎
@@
We can consider proper bimeromorphic morphisms between complex analytic orbispaces. They are closely related to proper bimeromorphic morphisms between the quotient spaces, as we explain as follows. Let $\mathfrak{X}=(X_{i},G_{i},\pi_{i})$ be a complex analytic orbispace with quotient space $X$. First we assume that there are $G_{i}$-equivariant proper bimeromorphic morphisms $f_{i}\colon Y_{i}\to X_{i}$ which are compatible along the overlap in the following sense. For any inclusions $\varphi_{i}\colon U\to X_{i}$ and $\varphi_{j}\colon U\to X_{j}$, the fiber products $U\times_{X_{i}}Y_{i}$ and $U{\times}_{X_{j}}Y_{j}$ are isomorphic over $U$. We note that such an isomorphism must be unique for $f_{i}$ and $f_{j}$ are bimeromorphic. Then the collection of the $(Y_{i},G_{i})$’s induces a complex analytic orbispace $\mathfrak{Y}$ with quotient space $Y$. Furthermore, there is an induced proper bimeromorphic morphism $f\colon Y\to X$, so that $X_{i}\times_{X}Y\cong Y_{i}$ and the corresponding basechange of $f$ coincides with $f_{i}\colon Y_{i}\to X_{i}$.
Conversely, we assume that there is a proper bimeromorphic morphism $f\colon Y\to X$ with $Y$ normal. Let $Y_{i}$ be the normalization of the main component of $X_{i}\times_{X}Y$. Then there is a natural action of $G_{i}$ on $Y_{i}$ and the natural projection $f_{i}\colon Y_{i}\to X_{i}$ is $G_{i}$-equivariant. Moreover, the collection of the $(Y_{i},G_{i})$’s defines a complex analytic orbispace with quotient space $Y$.
@@
Let $W=\mathbb{C}^{2}$ with coordinates $(a,b)$ and let $G=\mathbb{Z}/4\mathbb{Z}=<g>$. Then there is an action of $G$ on $W$ by setting $g.(a,b)=(-b,a)$. Then $(W,G,\pi)$ induces an orbifold with quotient space $X=W/G$. Let $D$ be the sum of the coordinates axes in $W$. Then $D$ is $G$-invariant and its image $E$ in $X$ is irreducible. We remark that the pair $(W,D)$ is snc, but $D$ does not induce a snc orbi-divisor, for $D=\pi^{-1}(E)$ is not smooth.
@@
Let $\mathfrak{X}=(X_{i},G_{i},\pi_{i})$ be a normal complex analytic orbispace with quotient space $X$. Assume that $\mathfrak{D}$ is an orbi-divisor in $\mathfrak{X}$. Then there are $G_{i}$-invariant projective bimeromorphic morphisms $f_{i}\colon Y_{i}\to X_{i}$, such that the collection of the $(Y_{i},G_{i})$’s induces a complex orbifold $\mathfrak{Y}$ with quotient space $Y$, and that the preimage of $\mathfrak{D}$ in $\mathfrak{Y}$ is a snc orbi-divisor.
@@
Let $E$ be the image of $\mathfrak{D}$ in $X$ and let $D_{i}=\pi_{i}^{-1}(E)$. Let $p_{i}\colon X^{\prime}_{i}\to X_{i}$ be the functorial principalization of the ideal sheaf of $D_{i}$, see Theorem 3.2 in the next section. Then up to replacing $X_{i}$ by $X^{\prime}_{i}$, we can assume that $X_{i}$ is smooth and $D_{i}$ is a snc divisor.
We note that, if $K$ and $K^{\prime}$ are strata of $D_{i}$ of minimal dimension, then either $K^{\prime}=K$ or $K^{\prime}\cap K=\emptyset$. Let $r$ be the minimum of the dimensions of the strata of $D_{i}$ for all $i$. Let $p_{i}\colon X^{\prime}_{i}\to X_{i}$ be the blowup at the union of all strata of $D_{i}$ of dimension $r$. Then the center of the blowup $p_{i}$ is either empty or smooth. Moreover, there is a natural action of $G_{i}$ on $X^{\prime}_{i}$ so that $p_{i}$ is $G_{i}$-invariant. The preimage $p_{i}^{-1}(D_{i})$ is a $G_{i}$-invariant snc divisor in $X_{i}^{\prime}$. Furthermore, the collection of the $(X^{\prime}_{i},G_{i})$’s induces a complex orbifold $\mathfrak{X}^{\prime}=(X^{\prime}_{i},G_{i},\mu_{i})$ with quotient space $X^{\prime}$. There is a natural projective bimeromorphic morphism $p\colon X^{\prime}\to X$, with a $p$-ample and $p$-exceptional divisor. If $\Delta\subseteq X^{\prime}$ is the exceptional set of $p$, then $\Delta$ is pure of codimension 1 and $\mu_{i}^{-1}(\Delta)$ is the exceptional set of $p_{i}$. In particular, $\mu_{i}^{-1}(\Delta)$ is smooth and is the disjoint union of its irreducible components.
We continue this procedure by blowing up $X^{\prime}_{i}$ at the union of all $(r+1)$-dimensional strata of the strict transform $(p_{i}^{-1})_{*}D_{i}$. By induction on the dimension of strata, we obtain projective bimeromorphic morphisms $f_{i}\colon Y_{i}\to X_{i}$ such that the following properties hold.
@@
    There is an induced bimeromorphic morphism $f\colon Y\to X$, whose exceptional locus is pure of codimension 1.
@@ 2.D. Minimal Model Programs for complex analytic spaces
Let $f\colon X\to Y$ be a projective surjective morphism between complex analytic spaces. We assume that either $Y$ is compact or $Y$ is a germ around a point $y\in Y$. Assume that $(X,\Delta)$ is a dlt pair and $X$ is $\mathbb{Q}$-factorial. We recall that $X$ is $\mathbb{Q}$-factorial if every Weil divisor in $X$ is $\mathbb{Q}$-Cartier, and some reflexive power $(\omega_{X}^{\otimes k})^{**}$ with $k\in\mathbb{Z}_{>0}$ of the canonical sheaf $\omega_{X}$ is locally free. In this case, we write $K_{X}$ for the class of $\frac{1}{k}c_{1}((\omega_{X}^{\otimes k})^{**})$. Then the $f$-relative cone theorem and contraction theorem hold, see [Nak87, Section 4], [DHP22, Section 2.6] and [Fuj22, Section 7]. Furthermore, if $H$ is a $f$-ample $\mathbb{Q}$-divisor, we can run a $f$-relative MMP for $(X,\Delta)$, with the scaling of $H$, see [DHP22, Theorem 1.4] or [Fuj22, Theorem 1.7]. The following lemma is essentially proved in [Bir12, Lemma 4.2].
@@
Let $X$ be a compact complex analytic variety. Let $\{S_{1},...,S_{r}\}$ be a collection of irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$ of codimension 2, such that $X$ has quotient singularities around general points of $S_{i}$ for all $i=1,...,r$. Then there is a projective bimeromorphic morphism $f\colon Y\to X$ such that the following properties hold.
@@
    $f$ is an isomorphism over general points of $S_{i}$ for all $i=1,...,r$.
@@
Let $r\colon\widetilde{X}\to X$ be a projective log resolution which is an isomorphism over $X_{\operatorname{\textsubscript{\rm sm}}}$. Let $E$ be the reduced sum of all $r$-exceptional prime divisors whose centers in $X$ are one of the $S_{i}$’s. Let $\varepsilon>0$ be a small enough rational number such that $-(1-\varepsilon)$ is smaller than the discrepancies for $K_{X}$ around general points of the $S_{i}$’s. Then, by [DHP22, Theorem 1.4], we can run a $r$-relative MMP for the klt pair $(\widetilde{X},(1-\epsilon)E)$, with the scaling of certain $r$-ample divisor, and obtain a projective bimeromorphic morphism $f\colon Y\to X$. Then $Y$ has $\mathbb{Q}$-factorial klt singularities. The negativity lemma shows that $E$ is contracted by $\widetilde{X}\dashrightarrow Y$. This implies that $f$ is an isomorphism over general points of every $S_{i}$.
@@
Let $(X,D)$ be a reduced pair, where $X$ is a normal compact complex analytic variety. Let $\{S_{1},...,S_{r}\}$ be the collection of all codimension 2 irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$, which are not contained in $D$. Assume that $X$ has quotient singularities around general points of $S_{i}$ for all $i=1,...,r$. Then there exists a projective bimeromorphic morphism $f\colon Y\to X$ such that the following properties hold.
@@
Let $\rho\colon Z\to X$ be a projective log resolution of $(X,D)$, which is an isomorphism over the snc locus of $(X,D)$. Let $E$ be the whole reduced $\rho$-exceptional divisor and let $\Gamma=E+\rho^{-1}_{*}D$. Then, thanks to [DHP22, Theorem 2.45], we can run a $\rho$-relative MMP for the dlt pair $(Z,\Gamma)$, with the scaling of some $\rho$-ample divisor $H$ as follows,
@@
If the MMP terminates, then we let $Y$ be the output of the MMP and let $f\colon Y\to X$ be the natural morphism. From the negativity lemma, we deduce that $f$ is an isomorphism over general points of each $S_{i}$.
Otherwise, by Lemma 2.17, the nef threshold sequence $(\lambda_{j})$ of the MMP tends to 0. Let $V_{i}$ be a local surface intersecting $S_{i}$ transversally at a very general point $x_{i}\in S_{i}$. Then the MMP induces a surface MMP over $V_{i}$. Since $X$ has quotient singularities around general points of $S_{i}$, we see that $V_{i}$ is a klt surface singularity by Lemma 2.1. Since $D$ does not contain $S_{i}$, we can pick $x_{i}\not\in D$. Then we deduce that there is some integer $N(i)\geq 0$ such that $Z_{j}$ is isomorphic to $X$ at $x_{i}$ if $j\geq N(i)$. Let $k$ be the maximum of $\{N(1),...,N(r)\}$. Then $Z_{k}\to X$ is an isomorphism over general points of every $S_{i}$. In this case, we set $Y=Z_{k}$ and denote by $f\colon Y\to X$ the natural morphism.

### v6

@@ 2.C. Notion of complex analytic orbispaces
    Let $U$ be a connected Hausdorff space. An orbifold chart of $U$ is a triple $(\widetilde{U},G,\pi)$ such that $\widetilde{U}$ is a reduced complex analytic space of dimension $n$, that $G$ is a finite group acting holomorphically on $\widetilde{U}$, and that $\pi\colon\widetilde{U}\rightarrow U$ is a continuous map inducing a homeomorphism from $\widetilde{U}/G$ to $U$. We denote by $\mathrm{ker}\,G$ the maximal subgroup of $G$ which acts trivially on $\widetilde{U}$.
@@
    Let $\iota\colon U^{\prime}\hookrightarrow U$ be an open embedding, and let $(\widetilde{U}^{\prime},G^{\prime},\pi^{\prime})$ be an orbifold chart of $U^{\prime}$. An injection from $(\widetilde{U}^{\prime},G^{\prime},\pi^{\prime})$ to $(\widetilde{U},G,\pi)$ consists of an open embedding $\varphi\colon\widetilde{U}^{\prime}\rightarrow\widetilde{U}$ and a group monomorphism $\lambda\colon G^{\prime}\rightarrow G$ such that $\varphi$ is equivariant with respect to $\lambda$, that $\lambda$ induces an isomorphism from $\mathrm{ker}\,G^{\prime}$ to $\mathrm{ker}\,G$, and that $\iota\circ\pi^{\prime}=\pi\circ\varphi$.
@@
Assume that $X$ is a complex analytic variety with klt singularities. Then for every point $x\in X$, there is an open neighborhood $U(x)$, and a smallest positive integer $k(x)$, such that $(\omega_{X}^{\otimes k(x)})^{**}|_{U(x)}\cong\mathcal{O}_{U(x)}$, where $\omega_{X}$ is the canonical sheaf of $X$. Let $V(x)\rightarrow U(x)$ be the cyclic cover induced by this isomorphism (see [KM98, Definition 2.52]), and let $G(x)$ be the Galois group. Then $V(x)$ has canonical singularities. Furthermore, the collection of the $(V(x),G(x))$’s induces a complex analytic orbispace with quotient space $X$, see [DO23, Lemma 2.3].
@@
Let $\mathfrak{X}=(X_{i},G_{i})_{i\in I}$ be a complex analytic orbispace with quotient space $X$. Let $\delta_{i}$ be a holomorphic function on $X_{i}$ such that the ideal sheaf $\mathcal{J}_{i}$ generated by $\delta_{i}$ is compatible along the overlaps. In other words, if $p_{i}\colon V\rightarrow X_{i}$ and $p_{j}\colon V\rightarrow X_{j}$ are inclusions of orbifold charts, then $p_{i}^{*}\mathcal{J}_{i}=p_{j}^{*}\mathcal{J}_{j}$ are the same ideal sheaf on $V$. Let $S\subseteq X$ be a compact subset and let $S_{i}=\pi_{i}^{-1}(S)$.
Then there is a finite family $(U_{k},G^{\prime}_{k})_{k\in K}$ of orbifold charts satisfying the following properties. There is an application $\sigma\colon K\rightarrow I$ and there are inclusions $\iota_{k}\colon U_{k}\rightarrow X_{\sigma(k)}$ of orbifold charts. Let $n>0$ be an integer and let $W_{k}=U_{k}[\sqrt[n]{\delta_{\sigma(k)}}]$. Then the natural morphism $W_{k}\rightarrow X$ is Galois over its image with Galois group $H_{k}$, and the family $(W_{k},H_{k})_{k\in K}$ induces a complex analytic orbispace, whose quotient space is an open neighborhood of $S$.
@@
For any $i\in I$, the assumption implies that there is a unit function $u_{g}$ such that $g^{*}\delta_{i}=u_{g}\cdot\delta_{i}$ for any $g\in G_{i}$. For every point $x\in S_{i}$, there is an open neighborhood $U_{x}$ of $x$ in $X_{i}$, such that $u_{g}|_{U_{x}}$ admits $n$-th roots. Let $G_{x}\subseteq G_{i}$ be the stabilizer of $x$. By shrinking $U_{x}$, we may assume that $(U_{x},G_{x})$ is an orbifold chart compatible with $\mathfrak{X}$. By the same argument of Lemma 2.5, the automorphisms of $U_{x}$ over $U_{x}/G_{x}$ can be lifted to automorphisms of $U_{x}[\sqrt[n]{\delta_{i}}]$. This implies that the natural morphism $U_{x}[\sqrt[n]{\delta_{i}}]\rightarrow X$ is Galois over its image.
By considering all points $x\in S_{i}$ and all $i\in I$, we obtain a family of orbifold charts $(U_{x},G_{x})$. Since $S$ is compact, we can extract a finite family of points $(x_{k})_{k\in K}$, with an application $\sigma\colon K\rightarrow I$, such that $x_{k}\in X_{\sigma(k)}$, and that the images of $U_{x_{k}}$ in $X$ covers $S$. We let $U_{k}=U_{x_{k}}$ and $G^{\prime}_{k}=G_{x_{k}}$. Then the previous paragraph implies that $W_{k}\rightarrow X$ is Galois over its image, with Galois group $H_{k}$. By Lemma 2.5 and Lemma 2.11, we deduce that the family $(W_{k},H_{k})_{k\in K}$ induces a complex analytic orbispace. ∎
@@
We can consider proper bimeromorphic morphisms between complex analytic orbispaces. They are closely related to proper bimeromorphic morphisms between the quotient spaces, as we explain as follows. Let $\mathfrak{X}=(X_{i},G_{i},\pi_{i})$ be a complex analytic orbispace with quotient space $X$. First we assume that there are $G_{i}$-equivariant proper bimeromorphic morphisms $f_{i}\colon Y_{i}\rightarrow X_{i}$ which are compatible along the overlap in the following sense. For any inclusions $\varphi_{i}\colon U\rightarrow X_{i}$ and $\varphi_{j}\colon U\rightarrow X_{j}$, the fiber products $U\times_{X_{i}}Y_{i}$ and $U{\times}_{X_{j}}Y_{j}$ are isomorphic over $U$. We note that such an isomorphism must be unique for $f_{i}$ and $f_{j}$ are bimeromorphic. Then the collection of the $(Y_{i},G_{i})$’s induces a complex analytic orbispace $\mathfrak{Y}$ with quotient space $Y$. Furthermore, there is an induced proper bimeromorphic morphism $f\colon Y\rightarrow X$, so that $X_{i}\times_{X}Y\cong Y_{i}$ and the corresponding basechange of $f$ coincides with $f_{i}\colon Y_{i}\rightarrow X_{i}$.
Conversely, we assume that there is a proper bimeromorphic morphism $f\colon Y\rightarrow X$ with $Y$ normal. Let $Y_{i}$ be the normalization of the main component of $X_{i}\times_{X}Y$. Then there is a natural action of $G_{i}$ on $Y_{i}$ and the natural projection $f_{i}\colon Y_{i}\rightarrow X_{i}$ is $G_{i}$-equivariant. Moreover, the collection of the $(Y_{i},G_{i})$’s defines a complex analytic orbispace with quotient space $Y$.
@@
Let $W=\mathbb{C}^{2}$ with coordinates $(a,b)$ and let $G=\mathbb{Z}/4\mathbb{Z}=<g>$. Then there is an action of $G$ on $W$ by setting g.(a,b)=(b,a)g.(a,b)=(-b,a). Then $(W,G,\pi)$ induces an orbifold with quotient space $X=W/G$. Let $D$ be the sum of the coordinates axes in $W$. Then $D$ is $G$-invariant and its image $E$ in $X$ is irreducible. We remark that the pair $(W,D)$ is snc, but $D$ does not induce a snc orbi-divisor, for $D=\pi^{-1}(E)$ is not smooth.
@@
Let $\mathfrak{X}=(X_{i},G_{i},\pi_{i})$ be a normal complex analytic orbispace with quotient space $X$. Assume that $\mathfrak{D}$ is an orbi-divisor in $\mathfrak{X}$. Then there are $G_{i}$-invariant projective bimeromorphic morphisms $f_{i}\colon Y_{i}\rightarrow X_{i}$, such that the collection of the $(Y_{i},G_{i})$’s induces a complex orbifold $\mathfrak{Y}$ with quotient space $Y$, and that the preimage of $\mathfrak{D}$ in $\mathfrak{Y}$ is a snc orbi-divisor.
@@
Let $E$ be the image of $\mathfrak{D}$ in $X$ and let $D_{i}=\pi_{i}^{-1}(E)$. Let $p_{i}\colon X^{\prime}_{i}\rightarrow X_{i}$ be the functorial principalization of the ideal sheaf of $D_{i}$, see Theorem 3.2 in the next section. Then up to replacing $X_{i}$ by $X^{\prime}_{i}$, we can assume that $X_{i}$ is smooth and $D_{i}$ is a snc divisor.
We note that, if $K$ and $K^{\prime}$ are strata of $D_{i}$ of minimal dimension, then either $K^{\prime}=K$ or $K^{\prime}\cap K=\emptyset$. Let $r$ be the minimum of the dimensions of the strata of $D_{i}$ for all $i$. Let $p_{i}\colon X^{\prime}_{i}\rightarrow X_{i}$ be the blowup at the union of all strata of $D_{i}$ of dimension $r$. Then the center of the blowup $p_{i}$ is either empty or smooth. Moreover, there is a natural action of $G_{i}$ on $X^{\prime}_{i}$ so that $p_{i}$ is $G_{i}$-invariant. The preimage $p_{i}^{-1}(D_{i})$ is a $G_{i}$-invariant snc divisor in $X_{i}^{\prime}$. Furthermore, the collection of the $(X^{\prime}_{i},G_{i})$’s induces a complex orbifold $\mathfrak{X}^{\prime}=(X^{\prime}_{i},G_{i},\mu_{i})$ with quotient space $X^{\prime}$. There is a natural projective bimeromorphic morphism $p\colon X^{\prime}\rightarrow X$, with a $p$-ample and $p$-exceptional divisor. If $\Delta\subseteq X^{\prime}$ is the exceptional set of $p$, then $\Delta$ is pure of codimension 1 and $\mu_{i}^{-1}(\Delta)$ is the exceptional set of $p_{i}$. In particular, $\mu_{i}^{-1}(\Delta)$ is smooth and is the disjoint union of its irreducible components.
We continue this procedure by blowing up $X^{\prime}_{i}$ at the union of all $(r+1)$-dimensional strata of the strict transform $(p_{i}^{-1})_{*}D_{i}$. By induction on the dimension of strata, we obtain projective bimeromorphic morphisms $f_{i}\colon Y_{i}\rightarrow X_{i}$ such that the following properties hold.
@@
    There is an induced bimeromorphic morphism $f\colon Y\rightarrow X$, whose exceptional locus is pure of codimension 1.
@@ 2.D. Minimal Model Programs for complex analytic spaces
Let $f\colon X\rightarrow Y$ be a projective surjective morphism between complex analytic spaces. We assume that either $Y$ is compact or $Y$ is a germ around a point $y\in Y$. Assume that $(X,\Delta)$ is a dlt pair and $X$ is $\mathbb{Q}$-factorial. We recall that $X$ is $\mathbb{Q}$-factorial if every Weil divisor in $X$ is $\mathbb{Q}$-Cartier, and some reflexive power $(\omega_{X}^{\otimes k})^{**}$ with $k\in\mathbb{Z}_{>0}$ of the canonical sheaf $\omega_{X}$ is locally free. In this case, we write $K_{X}$ for the class of $\frac{1}{k}c_{1}((\omega_{X}^{\otimes k})^{**})$. Then the $f$-relative cone theorem and contraction theorem hold, see [Nak87, Section 4], [DHP22, Section 2.6] and [Fuj22, Section 7]. Furthermore, if $H$ is a $f$-ample $\mathbb{Q}$-divisor, we can run a $f$-relative MMP for $(X,\Delta)$, with the scaling of $H$, see [DHP22, Theorem 1.4] or [Fuj22, Theorem 1.7]. The following lemma is essentially proved in [Bir12, Lemma 4.2].
@@
Let $X$ be a compact complex analytic variety. Let $\{S_{1},...,S_{r}\}$ be a collection of irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$ of codimension 2, such that $X$ has quotient singularities around general points of $S_{i}$ for all i=1,,ri=1,...,r. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ such that the following properties hold.
@@
    $f$ is an isomorphism over general points of $S_{i}$ for all i=1,,ri=1,...,r.
@@
Let $r\colon\widetilde{X}\rightarrow X$ be a projective log resolution which is an isomorphism over $X_{\operatorname{\textsubscript{\rm sm}}}$. Let $E$ be the reduced sum of all $r$-exceptional prime divisors whose centers in $X$ are one of the $S_{i}$’s. Let $\varepsilon>0$ be a small enough rational number such that $(1-\varepsilon)$ is smaller than the discrepancies for $K_{X}$ around general points of the $S_{i}$’s. Then, by [DHP22, Theorem 1.4], we can run a $r$-relative MMP for the klt pair $(\widetilde{X},(1-\epsilon)E)$, with the scaling of certain $r$-ample divisor, and obtain a projective bimeromorphic morphism $f\colon Y\rightarrow X$. Then $Y$ has $\mathbb{Q}$-factorial klt singularities. The negativity lemma shows that $E$ is contracted by $\widetilde{X}\dashrightarrow Y$. This implies that $f$ is an isomorphism over general points of every $S_{i}$.
@@
Let $(X,D)$ be a reduced pair, where $X$ is a normal compact complex analytic variety. Let $\{S_{1},...,S_{r}\}$ be the collection of all codimension 2 irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$, which are not contained in $D$. Assume that $X$ has quotient singularities around general points of $S_{i}$ for all i=1,,ri=1,...,r. Then there exists a projective bimeromorphic morphism $f\colon Y\rightarrow X$ such that the following properties hold.
@@
Let $\rho\colon Z\rightarrow X$ be a projective log resolution of $(X,D)$, which is an isomorphism over the snc locus of $(X,D)$. Let $E$ be the whole reduced $\rho$-exceptional divisor and let $\Gamma=E+\rho^{-1}_{*}D$. Then, thanks to [DHP22, Theorem 2.45], we can run a $\rho$-relative MMP for the dlt pair $(Z,\Gamma)$, with the scaling of some $\rho$-ample divisor $H$ as follows,
@@
If the MMP terminates, then we let $Y$ be the output of the MMP and let $f\colon Y\rightarrow X$ be the natural morphism. From the negativity lemma, we deduce that $f$ is an isomorphism over general points of each $S_{i}$.
Otherwise, by Lemma 2.17, the nef threshold sequence $(\lambda_{j})$ of the MMP tends to 0. Let $V_{i}$ be a local surface intersecting $S_{i}$ transversally at a very general point $x_{i}\in S_{i}$. Then the MMP induces a surface MMP over $V_{i}$. Since $X$ has quotient singularities around general points of $S_{i}$, we see that $V_{i}$ is a klt surface singularity by Lemma 2.1. Since $D$ does not contain $S_{i}$, we can pick $x_{i}\not\in D$. Then we deduce that there is some integer $N(i)\geq 0$ such that $Z_{j}$ is isomorphic to $X$ at $x_{i}$ if $j\geq N(i)$. Let $k$ be the maximum of $\{N(1),...,N(r)\}$. Then $Z_{k}\rightarrow X$ is an isomorphism over general points of every $S_{i}$. In this case, we set $Y=Z_{k}$ and denote by $f\colon Y\rightarrow X$ the natural morphism.